OpenStudy (anonymous):

The circle formed by the blades of a fan has a circumference of fifty-six and four-sevenths inches. What is its diameter? Use twenty-two sevenths for pi.

jimthompson5910 (jim_thompson5910):

Hint: The circumference of any circle is \[\Large C = \pi * d\] where C is the circumference and d is the diameter Since you're using \(\Large \frac{22}{7}\) for \(\Large \pi\), this means the formula becomes \[\Large C = \frac{22}{7}d\]